a bus travelling 40 mi/h and a car travelling 50 mi/h cover the same distance. The bus travels 1 hour more than the car. How many hours did each travel?
Answer by jim_thompson5910(35256)

d = rt

d = 40(t+1)

d = 50t

------------------------

40(t+1) = 50t

40t+40=50t

40t=50t-40

40t-50t=-40

-10t=-40

t=(-40)/(-10)

t=4

So the car traveled for 4 hours and the bus traveled for 5 hours.